# Navigating Scaling Laws: Compute Optimality in Adaptive Model Training

### v2

@@ 3 ViTs and Optimal Patch Sizes
Name
Width
Depth
Heads
Param (M)
GFLOPs
[…]
V256256256256-$6$/X
$256$
$6$
$8$
$5.1$
$1.22$
$0.120$
V192192192192-$12$/X
$192$
$12$
$3$
$5.6$
$1.43$
$0.136$
V256256256256-$12$/X
$256$
$12$
$4$
$9.9$
$2.44$
$0.240$
V384384384384-$12$/X
$384$
$12$
$6$
$21.8$
$5.25$
$0.538$
V512512512512-$12$/X
$512$
$12$
$8$
$38.6$
$9.13$
$0.953$
V640640640640-$12$/X
$640$
$12$
$10$
$60.0$
$14.1$
$1.49$
V768768768768-$12$/X
$768$
$12$
$12$
$86.2$
$20.1$
$2.14$
@@ 6.1 Adapting Model Width
[Uncaptioned image] Figure 10: Downstream performance as a function of compute for ViT of different size, trained with a patch size of 20. We use a log-log scale.
[Uncaptioned image] Figure 11: The theoretically predicted scheduled performance (left) and the empirical obtained (right) performance. While transitions are less smooth, the model based on the scheduler quickly recovers back to the predicted law.
@@ B.4 Scaling Laws
We evaluate on a held-out validation set. Although the validation samples are the same across different runs, we truncate them to match the context size of the respectively trained model. Models with longer contexts are thus expected to achieve lower perplexity due to the enhanced context size (conditioned on longer contexts, subsequent tokens are more predictable). As during inference we are interested in the validation perplexity over the longer contexts, we do not adjust for this discrepancy.
@@ Appendix D Adapting Multiple Model Shape Parameters Concurrently
We first present more results on which model configuration (number of parameters or patch size) leads to the most efficient training for different levels of performance in Fig. 19 and 20.
@@ Appendix G Discussion
Curriculum learning argues that the order in which samples are presented plays a crucial role in the learning efficiency of a model (Soviany et al., 2022). The role of training data undoubtedly played a crucial role in the convergence speed (Sorscher et al., 2022). Recently, other techniques for data selection have been proposed to accelerate large-scale pre-training (Evans et al., 2023). Our technique does not filter or select data, just chooses to invest different amounts of compute to different data, based on the current stage of training.

### v3

@@ 3 ViTs and Optimal Patch Sizes
Name Width Depth Heads Param (M) GFLOPs
$\textit{X}=8$ $\textit{X}=24$
V256256256256-$6$/X $256$ $6$ $8$ $5.1$ $1.22$ $0.120$
V192192192192-$12$/X $192$ $12$ $3$ $5.6$ $1.43$ $0.136$
V256256256256-$12$/X $256$ $12$ $4$ $9.9$ $2.44$ $0.240$
V384384384384-$12$/X $384$ $12$ $6$ $21.8$ $5.25$ $0.538$
V512512512512-$12$/X $512$ $12$ $8$ $38.6$ $9.13$ $0.953$
V640640640640-$12$/X $640$ $12$ $10$ $60.0$ $14.1$ $1.49$
V768768768768-$12$/X $768$ $12$ $12$ $86.2$ $20.1$ $2.14$
@@ 6.1 Adapting Model Width
Figure 10: Downstream performance as a function of compute for ViT of different size, trained with a patch size of 20. We use a log-log scale.
Figure 11: The theoretically predicted scheduled performance (left) and the empirical obtained (right) performance. While transitions are less smooth, the model based on the scheduler quickly recovers back to the predicted law.
@@ B.4 Scaling Laws
We evaluate on a held-out validation set. Although the validation samples are the same across different runs, we truncate them to match the context size of the respectively trained model. Models with longer contexts are thus expected to achieve lower test loss due to the enhanced context size (conditioned on longer contexts, subsequent tokens are more predictable). As during inference we are interested in the validation cross-entropy loss over the longer contexts, we do not adjust for this discrepancy.
@@
The exact number for the most compute-intensive point in Fig. 9 is $17825792000$ tokens. The exact model size including the embedding parameters is $137841408$ parameters. That leads to an approximate token per parameter value of $129.3$, past the optimal ”Chinchilla” point, leading thus to model that are more inference efficient.
@@ Appendix C Additional Experiments
We also include additional commonly used downstream performance metrics. In particular, we report 5/10-shot results on ImageNet/Pets/Birds as done in Zhai et al. (2022). Results can be seen in Fig. 18.
Figure 18: We include more downstream performance metrics for the V384384384384-$12$ model.
@@ Appendix D Adapting Multiple Model Shape Parameters Concurrently
We first present more results on which model configuration (number of parameters or patch size) leads to the most efficient training for different levels of performance in Fig. 20 and 21.
@@ Appendix G Discussion
Curriculum learning argues that the order in which samples are presented plays a crucial role in the learning efficiency of a model (Soviany et al., 2022). The role of training data undoubtedly played a crucial role in the convergence speed (Sorscher et al., 2022). Recently, other techniques for data selection have been proposed to accelerate large-scale pre-training (Evans et al., 2023). Our technique does not filter or select data, just chooses to invest different amounts of compute to different data, based on the current stage of training. Population based training is also a related area of work (Jaderberg et al., 2017).